# Bridging the micro-Hz gravitational wave gap via Doppler tracking with the Uranus Orbiter and Probe Mission: Massive black hole binaries, early universe signals and ultra-light dark matter

### v2

@@ IV.2 Model 2: Holodeck
We present another estimate of the detectable SMBHB population via the holodeck code [65, and see Ref. [2]]. Importantly, we utilise holodeck to build population models that generate nano-Hz-frequency GWBs with an amplitude that is consistent with the GWB for which multiple PTA experiments have recently found strong evidence [3, 6, 94, 119].
@@
To sample SMBHBs that are consistent with PTA observations, we generate 1080 holodeck populations by sampling over the uncertainties in each model parameter88829 parameters are marginalised over in this way, sampling from the PS__\__Astro__\__AStrong__\__AAll-v1.0 parameter space.. We then calculate the resulting PTA-measured GWB spectrum, fit a power-law spectrum to the lowest five frequency bins, and select the populations with amplitudes and spectral indices consistent with recent measurements. In particular, we require the GWB amplitude (characteristic strain at a frequency of $1\,\mathrm{yr}^{-1}$) that is $A_{\mathrm{yr}}\in[10^{-15},10^{-14}]$, and a power-law index (in power-spectral-density of timing residuals) that is $\gamma\in[-5.0,-2.5]$ [see, e.g., 109]. This procedure is repeated for each UOP sensitivity scenario and threshold SNR, drawing independent populations for each combination. Out of the 1080 populations, $130\pm 10$ pass the above selection cuts. Those populations are then integrated over with a cut in detection SNR similarly to Eq. (32) to determine detectable SMBHB systems.
@@ IV.3 Detection forecast
i.e., approximately linearly with a combination of the desired SNR and the achieved Allan deviation of the mission. Overall, the results shown in Table 1 suggest that a factor 30 improvement in the Allan deviation over the baseline mission would guarantee the detection of a few well constrained signals from SMBH binaries, in a mass range that is yet to be explored. Indeed, as seen in Fig. 8, the majority of detected sources would lie around a mass of few×$\times 10^{8}M_{\odot}$, being detectable out to redshifts of $z\sim 1$. The inspiral timescale of the typical source in the Millennium population estimate, i.e., $M\sim 10^{8}$ M, mass ratio of $\sim 1$ and frequency $f\sim 10^{-6}$ Hz, is approximately 10 yr. Therefore, the typical signals in the UOP are at the transition between almost monochromatic sources and chirping. They appear at frequencies of hundreds of nHz to tens of $\mu$Hz, i.e., above the PTA band and below the LISA band (see bottom panel of Fig. 8). A more sensitive mission can pick up more monochromatic signals at lower frequencies, where a larger number of binaries reside.
@@
Combining all of these prospects, we have demonstrated that Doppler tracking with the UOP is uniquely suited to detect GWs in the frequency regime between PTAs and space-based interferometry, and that Allan deviation improvements should be pursued and added to the mission’s technological requirements. Sources in this frequency range comprise heavy SMBH binaries of several $10^{8}M_{\odot}$ approaching coalescence, inspiraling pre-LISA systems and intermediate-mass black holes in the Milky Way. Such detections would unveil a more complete picture of SMBH formation, growth, and interaction, complementing both existing and planned GW detectors in the low-frequency regime. In the event of a concurrent detection with PTAs or LISA, a UOP detection may also aid in the difficult task of localising SMBHB sources [e.g., 106].
@@ VI Forecast constraints on ultralight dark matter
The expected constraints on $h_{c}$ from the top panel of Fig. 6 already give an idea of the reach of UOP to detect ULDM (compare Eqs. (1) and (8)). This suggests that for masses $m_{\scriptscriptstyle\rm DM}\gtrsim 10^{-22}$ eV/$c^{2}$, UOP may represent the most sensitive probe for large regions in parameter space of dark matter models, even in the baseline scenario. Notably, this mass range is particularly relevant for ULDM phenomenology [61, 43]. To illustrate this reach, let us consider the equivalent $h_{c}$ that would be produced by the three families of models we introduced in Section II.2, and that could be summarised by Eqs. (9), (10) and (12). Recall that $f_{\rm grav}\approx 2f_{\scriptscriptstyle\rm DM}$, where $f_{\scriptscriptstyle\rm DM}$ can be found in Eq. (6) for the purely gravitational coupling and the quadratic direct coupling $\alpha(\phi)=\phi^{2}/\Lambda^{2}$. For the linear coupling, such as the one in Eq. (11), $f=f_{\rm DM}$.
@@ VII.2 Breakthrough: Optical links?
[…]
@@ VIII Conclusion: Mission Outlook
  • The technological feasibility of reducing the Allan deviation with respect to its baseline value of $\sim 10^{-14}$ at 101010\,10s, as observed for the Cassini mission data more than three decades ago, in 1992.
@@ Appendix A White noise Bayesian sensitivity curve
Figure 11: We show the two-dimensional marginalised posterior distributions in amplitude and frequency for a set of three different injected GW amplitudes, corresponding to an SNR of 0, 3, and 12 from left to right. Here the SNR is calculated according to Eq. (20) with respect to the threshold amplitude shown in Fig. 12. As we can see, increasing SNR produces better parameter estimation in both frequency and amplitude of the GW, where sources with SNR=3 can already be constrained. The upper bound on the GW amplitude suffers from a degeneracy with the source’s position in the sky (not shown) and therefore stops improving significantly after an SNR8similar-toabsent8$\sim$ 8∼ 8 for this particular mission trajectory.
@@
To reiterate, the recovered $1\sigma$ threshold $A_{s}$, which corresponds to the definition of a sensitivity curve, results from distributing $N_{\rm t}$ tracking runs over the entire mission duration of $T=10$ years. For a monochromatic signal, this results in tracking a total of $n\sim(f_{\scriptscriptstyle\rm{GW}}\times N_{\rm t}\times 8\,\rm hour)$ cycles. This is reminiscent of matched filtering techniques, which enable one to compare the sensitivity curves of GW observatories to the characteristic strain of a signal, rather than its instantaneous amplitude [see, e.g., 96]. Thus, we adjust the $1\sigma$ detection threshold by heuristically factoring out the number of detected cycles:
@@
where $h_{\rm c}$ is the GW’s characteristic strain, and the factor 4 comes from the normalisation of the one-sided-power spectrum. Here, an SNR of $1$ corresponds to a barely detectable GW that can be distinguished from the absence of signal with $1\sigma$ confidence. As shown in Fig. 11, higher SNRs result in both better parameter estimation and in reduced false alarm rates, also in complete analogy with GW detectors. We find that the frequency and amplitude of sources with an SNR3similar-toabsent3$\sim$ 3∼ 3 can be recovered within $\sim 0.02$ and $\sim 0.5$ dex, respectively. The bounds improve by a factor $\sim 3$ for SNR $=12$ sources, though a residual degeneracy between the source’s amplitude and optimal localisation in the sky becomes manifest, inducing a skew in the recovered posterior distribution.

### v3

@@ IV.2 Model 2: Holodeck
We present another estimate of the detectable SMBHB population via the holodeck code [66, and see Ref. [2]]. Importantly, we utilise holodeck to build population models that generate nano-Hz-frequency GWBs with an amplitude that is consistent with the GWB for which multiple PTA experiments have recently found strong evidence [3, 6, 95, 120].
@@
To sample SMBHBs that are consistent with PTA observations, we generate 1080 holodeck populations by sampling over the uncertainties in each model parameter88829 parameters are marginalised over in this way, sampling from the PS_\_Astro_\_AStrong_\_AAll-v1.0 parameter space.. We then calculate the resulting PTA-measured GWB spectrum, fit a power-law spectrum to the lowest five frequency bins, and select the populations with amplitudes and spectral indices consistent with recent measurements. In particular, we require the GWB amplitude (characteristic strain at a frequency of $1\,\mathrm{yr}^{-1}$) that is $A_{\mathrm{yr}}\in[10^{-15},10^{-14}]$, and a power-law index (in power-spectral-density of timing residuals) that is $\gamma\in[-5.0,-2.5]$ [see, e.g., 110]. This procedure is repeated for each UOP sensitivity scenario and threshold SNR, drawing independent populations for each combination. Out of the 1080 populations, $130\pm 10$ pass the above selection cuts. Those populations are then integrated over with a cut in detection SNR similarly to Eq. (32) to determine detectable SMBHB systems.
@@ IV.3 Detection forecast
[…]
@@
Combining all of these prospects, we have demonstrated that Doppler tracking with the UOP is uniquely suited to detect GWs in the frequency regime between PTAs and space-based interferometry, and that Allan deviation improvements should be pursued and added to the mission’s technological requirements. Sources in this frequency range comprise heavy SMBH binaries of several $10^{8}M_{\odot}$ approaching coalescence, inspiraling pre-LISA systems and intermediate-mass black holes in the Milky Way. Such detections would unveil a more complete picture of SMBH formation, growth, and interaction, complementing both existing and planned GW detectors in the low-frequency regime. In the event of a concurrent detection with PTAs or LISA, a UOP detection may also aid in the difficult task of localising SMBHB sources [e.g., 107].
@@ VI Forecast constraints on ultralight dark matter
[…]
@@ VII.2 Breakthrough: Optical links?
For some orders of magnitude, let us put $\lambda=$1\text{\times}{10}^{-6}\text{\,}\mathrm{m}$\$, $A_{T}=A_{R}=$1\text{\,}\mathrm{m}^{2}$\$ and $R=$3\text{\times}{10}^{12}\text{\,}\mathrm{m}$\$. This makes the photon fraction $\sim 10^{-13}$ and the photon rate $\sim$5\text{\times}{10}^{4}\text{\,}\mathrm{W}^{-1}$\$.
With a passive reflector the photon fraction (37) gets squared through the round trip. At lunar distances ($R\simeq$4\text{\times}{10}^{8}\text{\,}\mathrm{m}$\$) this is tolerable, but at interplanetary distances, it would become hopeless. Hence an active system is needed.
@@ VIII Conclusion: Mission Outlook
  • The technological feasibility of reducing the Allan deviation with respect to its baseline value of $\sim 10^{-14}$ at 1010\,s, as observed for the Cassini mission data more than three decades ago, in 1992.
@@ Appendix A White noise Bayesian sensitivity curve
[…]
@@
To reiterate, the recovered $1\sigma$ threshold $A_{s}$, which corresponds to the definition of a sensitivity curve, results from distributing $N_{\rm t}$ tracking runs over the entire mission duration of $T=10$ years. For a monochromatic signal, this results in tracking a total of $n\sim(f_{\scriptscriptstyle\rm{GW}}\times N_{\rm t}\times 8\,\rm hour)$ cycles. This is reminiscent of matched filtering techniques, which enable one to compare the sensitivity curves of GW observatories to the characteristic strain of a signal, rather than its instantaneous amplitude [see, e.g., 97]. Thus, we adjust the $1\sigma$ detection threshold by heuristically factoring out the number of detected cycles:
@@
[…]
Appendix B Sky dependence
The typical sky-averaging of the GW strain is modified by the variable detector geometry of the UOP. We include a plot for visualisation purpuses in Fig. 13.
Figure 13: Sky dependence of the SNR for GW sources located at a given latitude and longitude with respect to the orbital plane and initial vector between the spacecraft and the Earth. Note the lack of a ”blind-spot” in the longitude, which is due to the curvature in the orbit.